OpenStudy (anonymous):

A sandbag was thrown downward from a building. The function f(t) = -16t2 - 64t + 80 shows the height f(t), in feet, of the sandbag after t seconds: Part A: Factor the function f(t) and use the factors to interpret the meaning of the x-intercept of the function. (4 points) Part B: Complete the square of the expression for f(x) to determine the vertex of the graph of f(x). Would this be a maximum or minimum on the graph? (4 points) Part C: Use your answer in part B to determine the axis of symmetry for f(x)? (2 points)

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

f(x) = -16x^2 - 64x + 512 A) f(x) = -16(x^2 + 4x - 32) = -16(x + 8)(x - 4) x-intercepts are x=-8, x=4 B) f(x) = -16(x^2 + 4x + 4) + 64 + 256 = -16(x + 2)^2 + 320 The vertex is (-2,320) C) The axis of symmetry is x=-2 2) 9x^2 - 16x + 60 = 0 A) b^2 - 4ac = 16^2 - 4*9*60 = -1904 There is no solution B) 4x^2 + 8x - 5 = 0 (2x + 5)(2x - 1) = 0 x = -5/2, x=1/2 I chose this method because it factored nicely. C) 2x^2 -12x + 5 = 0 x = [12 ± √(12^2 - 4*2*5)]/(2*2) = (12 ± √104)/4 = [12 ± √(4*26)]/4 = 3 ± √26/2 x = 0.4505 or x = 5.5495

OpenStudy (anonymous):

there ya go

OpenStudy (anonymous):

i got another question

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

wait where did you get +512 from

OpenStudy (anonymous):

ok i showed all my math

OpenStudy (anonymous):

f(x) = -16x^2 - 64x + 512 you put this as the equation

OpenStudy (anonymous):

yes i did

OpenStudy (anonymous):

but its f(t) = -16t2 - 64t + 80

OpenStudy (anonymous):

errrrr i tried

OpenStudy (anonymous):

lol its okay